alones (alones):

A piggy bank contains $4.45 is dimes and quarters. If the bank contains 25 coins, how many dimes are in the bank?

alones (alones):

i was thinking of13 but it was actually wrong :/

OpenStudy (123ab456c):

here i will gve you an example!:):):)

alones (alones):

sureee :)

OpenStudy (123ab456c):

Start by setting out what we know in equations. D = 0.10 Q = 0.25 X(0.10) + Y(0.25) = 2.25 X = 2Y Now, replace the X in the longer equation with its known Y equivalent. 2Y(0.10) + Y(0.25) = 2.25 Multiply out the brackets 0.2Y + 0.25Y = 2.25 Add the Y values together 0.45Y = 2.25 Divide both sides by 0.45 Y = 5 So, now we can replace the Y in the longer equation X(0.10) + 5(0.25) = 2.25 and resolve that X(0.10) + 1.25 = 2.25 X(0.10) = 2.25 - 1.25 X(0.10) = 1.00 X = 10 So, there are ten dimes and five quarters.

alones (alones):

O.O Wait the scond time i got 12 let me see

alones (alones):

thank you I GOT 12 AND IT WAS RIGHT!! THANKS YOU FOR THE EXAAMPLE<3

OpenStudy (123ab456c):

YOURWELCOME!

alones (alones):

=] Thanks

OpenStudy (123ab456c):

FAN AND MEDAL PLEASE!:)

alones (alones):

I think i medal you long ago hehe ;] i can re fan you :P

OpenStudy (123ab456c):

:):) GREAT JOB FOR ACTTUALY USING MY EXAMPLE ussually people like beg for the answer

alones (alones):

yeah thanks ;]

OpenStudy (123ab456c):

u deserve a medal

alones (alones):

THANK YOU<3 =]

OpenStudy (123ab456c):

NO THANK YOUR SELF YOU GOT IT RIGHT!!!:):):)GREAT JOB!!:):)

alones (alones):

hehe thanks ;]

OpenStudy (123ab456c):

;)